OpenStudy (anonymous):

find the value of y so that the line is passing through (9, 0) and (3, y) has a slope of -1/2

OpenStudy (anonymous):

\[Slope = \frac{y_2 - y_1}{x_2 - x_1}\]

OpenStudy (anonymous):

so how do i plug in the y?

OpenStudy (anonymous):

Put the given values in the formula: \[\huge \frac{-1}{2} = \frac{y - 0}{3-9}\] So, can you now solve it for y?

OpenStudy (anonymous):

so its -1/2 + y/6

OpenStudy (anonymous):

How?? Do it clearly??

OpenStudy (anonymous):

oh my bad i meant to put -1/2 = y/6

OpenStudy (anonymous):

-1/2 = y/-6

OpenStudy (anonymous):

Yes now you are right.. Multiply by (-6) both the sides and tell me what you get?

OpenStudy (anonymous):

6/-12 = y

OpenStudy (anonymous):

Solve it once again..

OpenStudy (anonymous):